OpenStudy (anonymous):

A mobile game uses two servers to handle matchmaking: one on the West Coast and one on the East Coast. The East Server has 50 free users and 20 premium, while the West server has 40 free users and 40 premium users. While the process for selecting the next user is random, there is a preference given for premium users: on each pick there is 60% than a premium user is selected and a 40% chance a free user is selected. What is the probability that the next player will be a free user from the East server?

OpenStudy (anonymous):

Can someone help me out here?

OpenStudy (anonymous):

Does anyone know how to this problem

OpenStudy (phi):

Prob of picking a free user from the east would be 50*0.4 / ( 60*0.6+90*0.4)

OpenStudy (anonymous):

How did you come up with those numbers

OpenStudy (phi):

they say selecting the next user is random, so I assume that means the pool of users they are choosing from is 150 user (70 from the east, 80 from the west) if the choices were equally weighted, the chance of a free user from the east would be 50/150 but the choices are not equal. the premiums are weight 0.6 and the free are weighted 0.4 that means break the 150 into 60 premium and 90 free and weight them to get 60*0.6 + 90*0.4 also, weight the 50 free up top, 50*0.4 we get 50*0.4 / ( 60*0.6+90*0.4) = 20/(36+36) = 20/72 = 5/18

OpenStudy (anonymous):

When I did the second part of the problem it came out as .16

OpenStudy (phi):

what is the second part of the problem ?

OpenStudy (anonymous):

Never mind I type in the wrong numbers

OpenStudy (anonymous):

Thanks for your help